# Data Trade and Consumer Privacy

### v2

@@ Abstract
This paper studies optimal mechanisms for collecting and trading data. Consumers benefit from revealing information about their tastes to a service provider because this improves the service. However, the information is also valuable to a third party as it may extract more revenue from the consumer in another market called the product market. The paper characterizes the constrained optimal mechanism for the service provider subject to incentive feasibility. It is shown that the service provider sometimes sells no information or only partial information in order to preserve profits in the service market. In a general setup, the service provision distortion and no-price discrimination in the product market are exclusive. Moreover, a ban on data trade may reduce social welfare because it makes it harder to price discriminate in the product market.
Keywords: Information market, Privacy, Price discrimination, Mechanism design, Information design
@@ 1 Introduction
The digital economy is deeply rooted in the sharing and use of personal data. Meanwhile, the transmission of personal data across markets brings concerns on privacy, [^1] For instance, personal data of Facebook users was collected by British consulting firm Cambridge Analytica in the 2010s. The privacy issues attract attention from both the public and regulatory authorities. The General Data Protection Regulation was put into effect in Europe in 2018, the California Consumer Privacy Act went into effect in 2020 and China’s Personal Information Protection Law was in force in late 2021. which can bite: The privacy concerns induced by data transmission, such as data trade, can adversely affect the data generating activities (e.g., data sharing, service usage) in related markets. The special feature of data endows it with a cross-market role.
Data is usually collected in one market and is also valuable in other markets. For example, Amazon Music, a music service provider, may collect taste data when users are enjoying the service of music recommendations. Amazon Music can trade the user data to a headphone seller (firm, producer, or data buyer) and the latter can use the data to price against users’ interests. As consumers have become increasingly aware of how their data can be collected, traded, and used, their willingness to use Amazon Music service and to share their data will be affected. With the cross-market role of data trade, the intermediary (service provider or platform) would design mechanisms for both collecting and trading data. What are the optimal mechanisms? How do the data trade and its cross-market effect shape the equilibrium privacy, surplus allocation, and social welfare?
To address these questions, this paper develops an integrated model including three sorts of interactions: (1) the intermediary’s data collection from consumers by providing service in the service market; (2) data trade between the intermediary and a firm in the data market; (3) pricing game between the firm and consumers in the product market. Figure 1 illustrates an overview of these interactions. [^2] The interactions in Figure 1 fits into many market scenarios such as Borrower-Fintech-Lender. In the example of Amazon Music, consumers share their taste information to obtain better music recommendations in the service market. Amazon Music sells the data to a headphone seller, who will sell headphones to consumers in the headphone market. We use the framework to study the intermediary optimal mechanisms of data collection and trade.
Our framework enables formal investigations on data sourcing and data trade in a comprehensive way. This contributes to the recent advocacy of integrated models to comprehensively understand the information markets (Bergemann and Bonatti, 2019) with the seminal envision back to Marschak (1968). Many works have separately investigated some specific elements, like data broker’s information selling mechanism or privacy protection, and less has offered a comprehensive investigation on data collection, data trade, and data usage. The holistic study of collection, trade, and usage of data and their effects and regulation implications cannot be easily discussed in divided models.
@@
In the framework, consumers who have private information about horizontal characteristics (tastes) benefit from sharing information with the intermediary by receiving a better-matched service but at the risk of the shared information being sold to a third party, the producer who can price against consumers using the bought data. In detail, consumers’ tastes are $H$ or $L$ (e.g., classical or rock music fans). Each consumer’s utility is from consuming service and product. The service provided to consumers may generate a mismatch as the intermediary does not know consumers’ true tastes. Each consumer also has unit demand for the product. The producer, a monopolist in the product market, does not know consumers’ willingness to pay for the product, which is related to consumers’ tastes. Hence, the taste data collected by the intermediary is valuable for the producer to set product prices. The intermediary can sell consumer data to the producer and it is the unique data source. The intermediary specifies (and commits to) a contract that consists of service provision, service fee, privacy policy (data trade), and data fee to consumers and the producer. Consumers and the producer decide whether to accept. Understanding that the data trade may discourage the service usage and data sharing of consumers, the intermediary’s contract would trade off between benefits of trading data and costs of hindering consumers’ data sharing and service usage.
@@
One key feature of the intermediary optimal mechanisms is that the optimal informativeness of data trade depends on comparisons of the horizontal differentiation level in the data-sourcing market (captured by the variation in service tastes) and vertical differentiation level in data-using market (captured by the variation in willingness to pay for product). The optimal data trade sells no or partial information when the horizontal differentiation level is low relative to the vertical one. In this case, as the service mismatch is small, consumers have less incentive to share their data to get the right service. The intermediary commits to selling less data to maintain consumers’ incentives to reveal information. Otherwise, suppose that the intermediary sells full information to the producer. Although the intermediary obtains most consumers surplus in the product market, it has to incentivize consumers’ information revelation by reducing service fees and hence loses the service profit. When the horizontal differentiation level is relatively low, the service profit losses outweigh the extra surplus gain in the product market. Hence, the intermediary would prefer to sell less data to preserve the service profits. In contrast, when the horizontal differentiation level is relatively high, the intermediary can sell full information. The service profit losses due to incentive distortions are small even if the intermediary sells full information. Under the full information data trade, consumers are price discriminated against in the product market.
The intermediary’s tradeoff holds in both the binary and general service space, but with different thresholds for the comparisons. As the service provision affects consumers’ incentives, we investigate a setup of general service in Section 5, in which the intermediary has more flexible service provisions and the service provided to the type-L consumers is downward distorted under some conditions. [^3] Section 4 assumes a binary service space to focus on the role of data trade, in which there is no service distortion to consumers in the intermediary optimal mechanisms. However, with these conditions, no-price discrimination in the product market cannot appear because the no-information data trade is not part of the optimal mechanisms. In other words, no-discrimination in product prices and distortion in the service provision are exclusive. In addition, in the optimal mechanisms, the intermediary extracts all the consumer surplus which is endogenous to data trade. This is because the data trade allows the intermediary to take back the consumer surplus of the product.
@@
Using the framework, we evaluate the effects of regulations on the data market. Our results suggest that the data trade ban does not improve social welfare. This happens because without any data it is difficult for the producer to price discriminate consumers. Suppose the producer believes that the product market consists of more type-H consumers, then without extra data the producer would set a high price and only sell to the type-Hs. Under the optimal data trade, the producer can extract the type-L consumers surplus by price discrimination, who are otherwise excluded from the market.
The findings have several implications. First, the data market itself can protect consumer privacy, for example, when the horizontal differentiation level in the data-sourcing market is low. Second, regulators need a holistic viewpoint over regulations on the data market. If regulators focus on only privacy protection or data trade, policies that aim at improving welfare may decrease it. Third, public monitoring of the intermediary’s implementation of data policy is needed. Otherwise, there can be a commitment issue with the intermediary’s privacy policy. The no-information data trade is optimal under some conditions from the cross-market perspective. But from an inter-temporal viewpoint, in practice, after the consumers reveal their data to the intermediary, it has incentives to trade consumers’ information for profits. Fourth, the co-existence of consumer data sharing and informative data trade in the optimal mechanisms rationalizes the so-called digital “privacy paradox”. [^4] It refers to the general discrepancies between consumers’ active data sharing behaviors and their self-stated privacy concerns. Several surveys, experimental, and empirical works are about this observation, see Gross and Acquisti (2005); Goldfarb and Tucker (2012); Athey et al. (2017); Chen et al. (2021) for example. The consumers’ information sharing and intermediary’s data selling as an equilibrium outcome are consistent with the recent evidence. Ichihashi (2023) attributes the paradox to consumers’ decreasing marginal privacy cost in a dynamic interaction with the platform.
@@
The paper is related to several strands of literature: information market, the economics of privacy, price discrimination, information design, and mechanism design.
@@
First, our paper adds to the literature on information markets (see recent surveys in Bergemann and Bonatti (2019) and Bergemann and Ottaviani (2021)). Admati and Pfleiderer (1990) focus on a monopolist that sells information about an asset and Bergemann and Bonatti (2015) study a data broker who sells cookies for targeting markets. Babaioff et al. (2012) and Bergemann et al. (2018) study optimal mechanisms for selling information in different contracting environments. Ichihashi (2021) and Bonatti et al. (2024) respectively investigate the competition among data intermediaries and data buyers. The majority of literature treats data sourcing as given. Few of these works investigate the mechanism designs for both collecting and selling information, which is the main attempt of the present paper. This paper highlights the cross-market effect of data trade.555 Consumer behaviors, especially as main data sources, can be distorted by the data trade. Argenziano and Bonatti (2023b) analyze strategic behaviors of privacy-conscious consumers. Several works show that the consumers’ manipulation incentives reduce the amount of information transmission in equilibrium, e.g., Frankel and Kartik (2022), Bonatti and Cisternas (2020).
Several works share a similar feature of tradeoffs between data exploitation and user privacy. Fainmesser et al. (2023) study the tradeoff between data collection and data protection but without a role of intermediary’s data trade, and hence the privacy is not endogenous to the data trade. [^6] They assume that the personal information (consumer activity such click) is verifiable. The privacy level is the protection investment. Argenziano and Bonatti (2023a) micro founds the privacy preferences. In our paper consumer privacy is based on the equilibrium data sale. Jullien et al. (2020) consider the tradeoff between data exploitation and user activity. Dosis and Sand-Zantman (2023) focus on a tradeoff between data monetization and processing. In these works, consumer information is verifiable. Our paper adopts the mechanism design approach because consumers control the information and information is unverifiable, which is more about consumers’ private characteristics than their verifiable activities, like browsing data. In addition, in our paper the tradeoff comes from the cross-market effect of data, relying on comparisons between the differentiation levels in data-sourcing and data-using markets.
Related to the cross-market effect of data, Doval and Skreta (Forthcoming) study the inter-temporal role of data with a focus on the tradeoff between product personalization and price discrimination in a dynamic model of upstream downstream. Rather, our paper investigates the cross-market role of data in a static environment. They find the no-price discrimination result under equal weight on profits of different periods in the commitment environment. No-discrimination can show up in the product market only if there is no distortion in service provisions to consumers in the data sourcing market.
This paper also belongs to the literature on price discrimination. Bergemann et al. (2015) study how market segmentation can support an arbitrary surplus distribution between consumers and a monopolist. The endogenous market segment also arises from consumers’ equilibrium disclosure about their characteristics in Ichihashi (2020) and Ali et al. (2023) and data broker’s creation and sale of market segments for profits in Yang (2022). This paper complements the literature by focusing on the critical role of the intermediary in both collecting data and trading data.
@@
The price discrimination endogenous to the information is related to the economics of privacy with seminal works of Taylor (2004) and Calzolari and Pavan (2006); see Acquisti et al. (2016) and Goldfarb and Que (2023) for recent surveys. The privacy concerns arise from sellers’ learning about consumers’ type (Villas-Boas, 2004; Conitzer et al., 2012) and buyers’ information revelation to sellers (Hidir and Vellodi, 2021; Ichihashi, 2020). Argenziano and Bonatti (2023a) study the data linkages instead of the intermediary’s data selling. Ali et al. (2023) combine the consumer information disclosure and personalized pricing. In dynamic environments, the privacy concerns on discrimination induce a ratchet effect (Laffont and Tirole, 1988). The ratchet effect can bite all the information value as the data collecting intermediary gets better off committing to offering full privacy (Calzolari and Pavan, 2006). Doval and Skreta (Forthcoming) find in a limited commitment environment that the price discrimination induces product line gap due to this effect. Rhodes and Zhou (2021) study consumer privacy choice in the context of personalized pricing. Our paper extends the literature by endogenizing the privacy policy as a part of the intermediary’s holistic design of information collection and sale.
In addition, it is related to the broad literature of information design (Kamenica and Gentzkow, 2011; Bergemann and Morris, 2019) and mechanism design. [^7] See Mussa and Rosen (1978), Myerson (1981) for classic works and Bergemann and Välimäki (2006) for a survey on information in mechanism design. In our paper consumers, as the data sources, have private unverifiable information. Hence, we rely on the mechanism design approach to incentivize consumers’ information revelation and the privacy policy or data trade is one of the intermediary’s design elements. Bergemann et al. (2018) study the mechanism of selling information. Bonatti et al. (2024) investigate data selling to competing buyers and Doval and Skreta (Forthcoming) study personalization of product and prices also in a mechanism design framework. We embed the information design approach to selling information into the intermediary’s mechanism design problem of incentivizing consumers’ information revelation to highlight the cross-market role of data trade. [^8] Yoder (2022) explores the contract to incentivize agent’s information acquisition and Georgiadis and Szentes (2020) study the designs of wage and performance evaluation policies as incentive tools. In our paper, one feature of optimal mechanisms is that the intermediary, as the designer, can extract all consumers surplus as the consumers surplus is endogenous.
Lastly, although this paper abstracts from data externalities, the finding that data trade improves social welfare is related to the works studying externality in the data market. Acemoglu et al. (2022) find that banning the data market improves welfare as the over-sharing of data in the market induced by data externality disappears. The welfare loss comes from the firm’s over-collection of consumer data in Bergemann et al. (2022) and Choi et al. (2019).
@@
The rest of this paper is organized as follows. Section 2 sets up the model. Section 3 takes some preliminary analyses of individual decisions. Section 4 characterizes the intermediary optimal mechanism under binary service and discusses regulations of the data market. Section 5 extends to a general service space. Section 6 presents some discussions and section 7 concludes. Omitted proofs and additional results are relegated to Appendices.
2 Model

### v3

@@ Abstract
This paper studies optimal mechanisms for collecting and trading data. Consumers benefit from revealing information about their tastes to a service provider because this improves the service. However, the information is also valuable to a third party as it may extract more revenue from the consumer in another market called the product market. The paper characterizes the constrained optimal mechanism for the service provider subject to incentive feasibility. It is shown that although the service provider sometimes sells no information or only partial information in order to preserve profits in the service market, selling full information is optimal when the data-sourcing market is highly differentiated. Moreover, a ban on data trade may reduce social welfare because it makes it harder to price discriminate in the product market. Instead, reducing the intermediary’s bargaining power can protect privacy without hurting social welfare, which suggests that the regulation of market power is more efficient than the regulation of data sharing.
Keywords: Data market and regulation, Intermediary, Privacy, Price discrimination, Mechanism design, Information design
@@ 1 Introduction
The digital economy is deeply rooted in the sharing and use of personal data. Meanwhile, the transmission of personal data across markets brings concerns about privacy, [^1] For instance, personal data of Facebook users was collected by British consulting firm Cambridge Analytica in the 2010s. The privacy issues attract attention from both the public and regulatory authorities. The General Data Protection Regulation (GDPR) was put into effect in Europe in 2018, the California Consumer Privacy Act (CCPA) went into effect in 2020 and China’s Personal Information Protection Law was in force in late 2021. which can bite: The privacy concerns induced by data transmission, such as data trade, can adversely affect the data generating activities (e.g., data sharing, service usage) in related markets. The special feature of data endows it with a cross-market role.
Data is usually collected in one market and is also valuable in other markets. For example, Amazon Music, a music service provider, may collect taste data when users are enjoying the service of music recommendations. Amazon Music can trade the user data to a headphone seller (firm, producer, or data buyer), and the latter can use the data to price against users’ interests. As consumers have become increasingly aware of how their data can be collected, traded, and used, their willingness to use Amazon Music service and to share their data will be affected. With the cross-market role of data trade, the intermediary (service provider or platform) would design mechanisms for both collecting and trading data. What are the optimal mechanisms? How do the data trade and its cross-market effect shape the equilibrium privacy, surplus allocation, and social welfare? Does the regulation of data sharing work better than the regulation of market power?
To address these questions, this paper develops an integrated model including three sorts of interactions: (1) the intermediary’s data collection from consumers by providing service in the service market; (2) data trade between the intermediary and a firm in the data market; (3) pricing game between the firm and consumers in the product market. Figure 1 illustrates an overview of these interactions. We use the framework to study the intermediary optimal mechanisms of data collection and trade, and data market regulations.
@@
Our framework enables formal investigations on data sourcing and trade in a comprehensive way, and hence the equilibrium interactions between second degree price discrimination in the data-sourcing market and third degree discrimination endogenous to data trade. Many works have separately investigated some specific elements, like information collection by screening (Mussa and Rosen, 1978), information sales by certification (Lizzeri, 1999), or data brokers’ information selling mechanism (Yang, 2022), and less has offered a comprehensive investigation on data collection, data trade, and data usage. The holistic study of the intermediary’s collection, trade, and usage of data and their effects and regulatory implications cannot be easily discussed in divided models. This framework contributes to the advocacy of integrated models to comprehensively understand the information markets (Bergemann and Bonatti, 2019) with the seminal envision back to Marschak (1968).
The framework also offers to study the intermediary’s bargaining power. The distribution of bargaining power is important when it comes to the transfer of data across markets via intermediation, because transferring the data from consumers to the intermediary does not necessarily transfer the bargaining power to the intermediary. Although a large literature has looked at the comparative statics across the information available to the market participants, [^2] E.g., the information available to the seller (Bergemann et al., 2015) or to the buyer (Roesler and Szentes, 2017) or to both (Kartik and Zhong, 2023). the information intermediary’s bargaining power and its regulation implications are under-explored. [^3] See discussions about the bargaining power of a durable good reseller in Calzolari and Pavan (2006a). We discuss the literature more fully below.
In the framework of Figure 1, consumers who have private information about horizontal characteristics (tastes) benefit from sharing information with the intermediary by receiving a better-matched service but at the risk of the shared information being sold to a third party, the producer who can price against consumers using the bought data. In detail, consumers’ tastes are $H$ or $L$ (e.g., classical or rock music fans). Each consumer’s utility is from consuming the service and product. The service provided to consumers may generate a mismatch as the intermediary does not know consumers’ true tastes. Each consumer also has unit demand for the product. The producer, a monopolist in the product market, does not know consumers’ willingness to pay for the product, which is related to consumers’ tastes. Hence, the taste data collected by the intermediary is valuable for the producer to set product prices. The intermediary can sell consumer data to the producer and it is the unique data source. The intermediary specifies (and commits to) a contract that consists of service provision, service fee, privacy policy (data trade), and data fee to consumers and the producer simultaneously. Consumers and the producer decide whether to accept. Understanding that the data trade may discourage the service usage and data revelation of consumers, the intermediary’s contract would trade off between benefits of trading data and costs of hindering consumers’ data sharing and service usage. [^4] The interactions in Figure 1 fit into many market scenarios such as Borrower-Fintech-Lender, in which consumers reveal their personal data while using the financial service provided by banks. Banks may transfer data via Application Programming Interface, i.e., open banking, to third-party fintech firms, which may offer personalized interest rates to consumers as borrowers in the lending market. In the example of Amazon Music, although how much a consumer listens to music could also be indicative of willingness to pay for headphones, the music service provider usually can verify how much a consumer listens to from the history, which is more in the spirit of the behavior-based price discrimination (see, e.g., Fudenberg and Villas-Boas (2006)).
@@
One key feature of the intermediary optimal mechanisms is that the optimal informativeness of data trade depends on comparisons of the horizontal differentiation level in the data-sourcing market (captured by the variation in service tastes) and vertical differentiation level in the data-using market (captured by the variation in willingness to pay for the product). The optimal data trade sells no or partial information when the horizontal differentiation level is sufficiently low relative to the vertical one. [^5] Although the intermediary protects privacy in Calzolari and Pavan (2006b), the setups and driving forces are different, as discussed in the literature below. In this case, as the service mismatch is small, consumers have less incentive to share their data to get the right service. The intermediary commits to selling less data to maintain consumers’ incentives to reveal information. Otherwise, suppose that the intermediary sells full information to the producer. Although the intermediary obtains most consumers surplus in the product market, it has to incentivize consumers’ information revelation by reducing service fees and hence loses the service profit. When the horizontal differentiation level is relatively low, the service profit losses outweigh the extra surplus gain in the product market. Hence, the intermediary would prefer to sell less data to preserve the service profits. In contrast, when the horizontal differentiation level is high enough, the intermediary can sell full information. The service profit losses due to incentive distortions are small even if the intermediary sells full information.
The second remarkable result is that introducing the intermediary’s partial bargaining power in the data market can decrease the amount of information disclosed by the optimal mechanism, but does not affect the surpluses of the intermediary, consumers, and producer. First, decreasing informativeness has two effects: the service profit gain and the data profit loss. With the intermediary’s full bargaining power, these two effects cancel each other out, leading to multiple solutions of disclosure policy; with partial bargaining power, the service profit gain outweighs the data profit loss, which drives the multiple solutions of data trade to the unique one, i.e., the minimum information among them. Second, the partial bargaining power over data profit does not affect the intermediary’s payoff because the equilibrium data profit is zero, while the consumer and producer get their outside option.
@@
Using the framework, we study regulations on the data market. In particular, we focus on regulations of the intermediary’s data sharing and market power. First, our results show that the data trade ban does not improve social welfare. This happens because without any data it is difficult for the producer to price discriminate among consumers. Under the optimal data trade, the producer can extract by price discrimination the surpluses of all the consumers, some of whom are otherwise excluded from the market. Second, regulating the intermediary’s power (how much the intermediary can extract the data revenue) will protect consumer privacy without hurting social welfare, because reducing the intermediary’s bargaining power doesn’t decrease the surpluses nor total welfare. The findings suggest that the regulation of the intermediary’s market power would be more efficient than the regulation of the intermediary’s data sharing.
@@
The paper relates to several strands of literature: information market, the economics of privacy, price discrimination, information design, and mechanism design.
@@
First, our paper adds to the literature on information markets (see recent surveys in Bergemann and Bonatti (2019) and Bergemann and Ottaviani (2021)). Admati and Pfleiderer (1990) focus on a monopolist that sells information about an asset, and Lizzeri (1999) focuses on an informed seller who pays for the intermediary’s certification to signal quality. Bergemann and Bonatti (2015) study a data broker who sells cookies for targeting markets; Babaioff et al. (2012) and Bergemann et al. (2018) study optimal mechanisms for selling information in different contracting environments. Ichihashi (2021) and Bonatti et al. (2024) respectively investigate the competition among data intermediaries and data buyers. Park and Evans (2025) study the intermediary’s incentives of information design to influence the consumer’s willingness to pay for information and the seller’s pricing decision, with a focus on the consumer’s purchase of information. The majority of literature treats data sourcing as given; few of them investigate the intermediary’s collection and sale of information. [^6] Consumer behaviors, especially as main data sources, can be distorted by the data trade. Argenziano and Bonatti (2023) analyze strategic behaviors of privacy-conscious consumers. Several works show that the consumers’ manipulation incentives reduce the amount of information transmission in equilibrium, e.g., Frankel and Kartik (2022), Bonatti and Cisternas (2020).
One notable exception is Calzolari and Pavan (2006b), that study disclosure in a sequential contracting setting and find that the upstream principal offers full privacy for the agent. Our simultaneous contracting setting admits selling information to be optimal. Their full privacy result requires the intermediary’s preference to have the same sign of the single-crossing condition in terms of trade or not with the third party, but the preference structure has different signs in our paper. The insight in the simultaneous setting is driven by the interaction between third degree price discrimination in the data-using market and second degree price discrimination in the data-sourcing market. Moreover, Calzolari and Pavan (2006b) do not consider the bargaining power of the intermediary, and Calzolari and Pavan (2006a) discuss the bargaining power but don’t study its regulatory implications. Our paper highlights both the data sourcing and the intermediary’s bargaining power.
@@
Second, the paper is related to the economics of privacy with seminal works of Taylor (2004) and Calzolari and Pavan (2006b); see Acquisti et al. (2016) and Goldfarb and Que (2023) for recent surveys. The privacy concerns arise from sellers’ learning about consumers’ type (Villas-Boas, 2004; Conitzer et al., 2012) and buyers’ information revelation to sellers (Hidir and Vellodi, 2021; Ichihashi, 2020). Argenziano and Bonatti (Forthcoming) study the data linkages instead of the intermediary’s data selling. Ali et al. (2023) combine the consumer information disclosure and personalized pricing. In dynamic environments, the privacy concerns on discrimination induce a ratchet effect (Laffont and Tirole, 1988). The ratchet effect can bite all the information value as the data collecting intermediary gets better off committing to offering full privacy (Calzolari and Pavan, 2006b). Doval and Skreta (2025) find in a limited commitment environment that the price discrimination induces a product line gap due to this effect. Rhodes and Zhou (2021) study consumer privacy choice in the context of personalized pricing. Our paper extends the literature by endogenizing the privacy policy as a part of the intermediary’s design of information collection and sale.
Several other works share a similar feature of tradeoffs between data exploitation and user privacy. Fainmesser et al. (2023) study the tradeoff between data collection and data protection but without a role of intermediary’s data trade, and hence the privacy is not endogenous to the data trade. [^7] They assume that the personal information (consumer activity such as clicks) is verifiable. The privacy level is the protection investment. Argenziano and Bonatti (Forthcoming) micro founds the privacy preferences. In our paper consumer privacy is based on the equilibrium data sale. Jullien et al. (2020) consider the tradeoff between data exploitation and user activity. Dosis and Sand-Zantman (2023) focus on a tradeoff between data monetization and processing. In these works, consumer information is verifiable. Our paper adopts the mechanism design approach because consumers control the information and information is unverifiable, which is more about consumers’ private characteristics than their verifiable activities, like browsing data. Related to the cross-market role of data, Doval and Skreta (2025) study the inter-temporal role of data. They focus on the tradeoff between product personalization and price discrimination in a dynamic model. Rather, our paper investigates in a static environment the intermediary’s balance between the data sourcing and data sales and how the intermediary’s bargaining power affects the balance.
Third, this paper also adds to the literature on price discrimination and market segmentation by emphasizing the endogenous interactions between the second degree and third degree price discrimination. Bergemann et al. (2015) study how market segmentation can support an arbitrary surplus distribution between consumers and a monopolist. [^8] Haghpanah and Siegel (2022) identify markets for which the entire surplus triangle of Bergemann et al. (2015) is achievable and Haghpanah and Siegel (2023) study Pareto improving market segmentation with multiple products. The endogenous market segment also arises from consumers’ equilibrium disclosure about their characteristics in Ichihashi (2020) and Ali et al. (2023) and data brokers’ creation and sale of market segments for profits in Yang (2022). So our framework contributes to discussing the relation between the third degree price discrimination by market segmentation and the second degree discrimination induced by the monopoly screening.
Lastly, it is related to the broad literature of information design (Kamenica and Gentzkow, 2011; Bergemann and Morris, 2019) and mechanism design. [^9] See, e.g., Mussa and Rosen (1978), Myerson (1981) for classic works and Bergemann and Välimäki (2006) for a survey on information in mechanism design. Several works study the comparative statics of information available to the seller (Bergemann et al., 2015) or to the buyer (Roesler and Szentes, 2017) or to both (Kartik and Zhong, 2023). Bergemann et al. (2018) study the mechanism of selling information. Yang (2022) studies a setting where the data buyer has private information about costs. Bonatti et al. (2024) and Doval and Skreta (2025) investigate mechanism design problems with multiple buyers and multiple periods, respectively. They don’t consider the source of information. In our paper, consumers, as the data sources, have private unverifiable information. Hence, we rely on the mechanism design approach to incentivize consumers’ information revelation, and the privacy policy or data trade is one of the intermediary’s design elements. The study of the interaction between monopoly screening and endogenous market segmentation complements the standard monopoly screening (Mussa and Rosen, 1978; Maskin and Riley, 1984), which does not consider the information disclosure as part of allocations. Calzolari and Pavan (2006b) study screening plus persuasion and Dworczak (2020) confines to cut-off mechanism in sequential contracting. This paper differs in the simultaneous setup and complements the literature by looking at the critical role of the intermediary’s bargaining power.
@@
The rest of this paper is organized as follows. Section 2 sets up the model. Section 3 takes some preliminary analyses of individual decisions. Section 4 investigates the intermediary optimal design: Section 4.1 focuses on the role of data trade, Section 4.2 completely characterizes the intermediary optimal mechanism and discusses the regulation of data sharing. Section 5 studies the intermediary’s bargaining power in the data market and the regulation of the intermediary’s power. Section 6 provides some discussions and Section 7 concludes. Omitted proofs and additional results are relegated to Appendices.
2 The Model